OpenStudy (anonymous):

sinxcosy = 1/2 (sin(x+y) + cos(x-y)) True or False?

OpenStudy (anonymous):

sin(x+y) = sin(x)cos(y) + sin(y)cos(x) cos(x-y) = cos(x)cos(y) + sin(x)sin(y) can do you it now ?

OpenStudy (anonymous):

er true?

OpenStudy (anonymous):

\[\sin(x+y) = \sin(x) \cos(y) + \cos(x) \sin(y)\] \[\cos(x-y) = \cos(x) \cos(y) + \sin(x) \sin(y)\]

OpenStudy (anonymous):

Can you add now \(sin(x+y)\) and \(cos(x-y)\) ??

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

its (sin(x+y)) + (cos(x-y)) ._.

OpenStudy (anonymous):

I said add expansion that we have given there above.. Add the expansion of sin(x+y) and cos(x-y) and tell me what did you get ??

OpenStudy (anonymous):

i just told you what "I" can get

OpenStudy (anonymous):

using : sin(x+y) = sin(x)cos(y) + sin(y)cos(x) cos(x-y) = cos(x)cos(y) + sin(x)sin(y) your equation becomes sinxcosy = 1/2 (sin(x)cos(y) + sin(y)cos(x) + cos(x)cos(y) + sin(x)sin(y))

OpenStudy (anonymous):

so ....true then

OpenStudy (anonymous):

no false bc thats sin x cos y = sinx cosy ...wait ..waht?

OpenStudy (anonymous):

show me

OpenStudy (calculusfunctions):

Wow! @Mpost knows how to show off but is neither elegant in his writing nor can he solve such simple simple questions. It seems like the only thing he does know is how to be disrespectful towards authority figures. Needs serious counselling. Almost feel sorry for this arrogant sorry excuse for a human.

OpenStudy (anonymous):

It is wrong...

OpenStudy (anonymous):

\[\rm 2\sin(x)\cos(y) \ne \sin(x)\cos(y) + \sin(y)\cos(x) + \sin(x)\sin(y) + \cos(x)\cos(y)\]